You purchased 1,400 shares of Barrett Golf Corp. stock at a price of $36.36 per share. While you owned the stock, you received dividends totaling $.67 per share. Today, you sold your stock at a price of $40.31 per share. What was your total dollar return on the investment?
Answer #1

Calculation of return:

Return per share= Selling price-Purchase price+ dividend

Return per share= 40.31-36.36+0.67= 4.62

Dollar return= 1400*4.62= 6468

Dollar return= $6468